# Dynamic LLM-Agent Network: An LLM-agent Collaboration Framework with Agent Team Optimization

### v1

@@ D.1 Human Priors and Automatic Evaluation Results
We further investigated how these agents selected by our unsupervised metric Agent Importance Score differ from human priors (e.g., these predefined roles). To do so, we calculated Agent Importance Scores for 7 agents on each subject of the MMLU dataset. As an example, we show the subjects where the agent of ”Doctor” and ”Programmer” has the highest agent importance score among all agents in Table 7 and Table 8. Though most subjects seems to be reasonably aligned with the role of the agent based on human priors (with green annotations), there are some subjects that do not match human priors, e.g., high school computer science as the subject that “Doctor” has the highest score. It exhibits the difference between human priors and the evaluation results Agent Importance Scores on other agents.
@@
D.2 Stability of DyLAN on Temperature
Figure 4: Performance of different methods under low and high temperatures on MATH (left) and HumanEval (right) datasets. DyLAN shows better robustness to different temperature and even takes advantage of higher temperature.
We tested a few methods on MATH (with simple CoT prompts) and HumanEval datasets under both low and high temperatures and repeated each experiment 3 times when the temperature was not 0. We exhibit the experimental results in Figure 4. From experimental results, we found that DyLAN is more stable on different hyper-parameters.
Experiments show that temperature greatly influences arithmetic reasoning and code generation tasks. In Figure 4, we found that most baseline methods have significant performance drops when temperature increases, but DyLAN shows strong robustness to various temperatures. We surprisingly found that DyLAN gets better results when temperature rises, suggesting it has benefited from diversity instead of being disturbed by low-quality answers of high-temperature agents. The inference-time agent selection may lead to the higher accuracy by keeping best responses when agents’ replies become more diverse. In conclusion, the collaboration of different roles functions effectively and robustly in the dynamic architecture. Nonetheless, higher temperature requires DyLAN to take more API calls (about +0.98 on average on MATH (temperature: $0.2\rightarrow 1.0$)).
D.3 Different Ranking Methods
Table 9: Overall accuracy (%) of DyLAN with different ranking method in the inference-time agent selection on MMLU dataset. Other settings are identical with Table 3.
@@ D.3 Different Ranking Methods
We also tested different ranking methods for inference-time agent selection of DyLAN on the MMLU dataset. We tested listwise ranker with our own prompts, pairwise GPT ranker from original LLM-Blender (Jiang et al., 2023), Elo Score from TrueSkill (Herbrich et al., 2006) also implemented with pairwise ranker, and pairwise ranker with Sliding Window algorithm (Qin et al., 2023). In Table 9, we show that different ranking methods have a relatively low impact on performance, but pairwise ranking methods always consume higher computational cost. Thus, we chose a listwise ranker in our implementation of DyLAN.
D.4 Relation between Domains and Optimized Teams of Agents
In the agent team optimization experiments on the DyLAN of 7 agents with different roles, we recorded the in-domain rate with different numbers of remaining roles in Table 10. The in-domain rate is the proportion of teams with at least one agent matching the domain of the subject based on human priors (e.g. “Programmer” matches STEM subjects, as described in Section 4.1), given that the teams are optimized on each subject. The results suggest that the optimized team of agents is roughly aligned with human priors but still with a significant offset.
Table 10: In-domain rate of teams that contain roles in the domain of the subject after optimization on the DyLAN of 7 agents with distinct roles.
#Roles 7 6 5 4 3 2 1
In-Domain Rate 1.00 0.97 0.91 0.88 0.74 0.60 0.39
D.5 Shapley Value v.s. Agent Importance Score
@@ D.4 Relation between Domains and Optimized Teams of Agents
Shapley Value is a widely used supervised metric for evaluating contribution of a single agent in a multi-agent system. We implement a simplified algorithm for LLM-agent collaboration systems. Given that the collaboration process is symmetric to roles in the formulation of the feed-forward network, we could reduce the permutation set in the original formula (Lundberg & Lee, 2017) to the combination set:
$S_{i}({\mathbb{R}})=\frac{1}{|{\mathbb{C}}||{\mathbb{R}}|}\sum_{{\mathbb{T}}\in{\mathbb{C}}}\frac{}{}(\mathrm{Performance}({\mathbb{T}}\cup\{i\})-\mathrm{Performance}({\mathbb{T}})),$ (2)
@@
where ${\mathbb{R}}$ is the set of agents in the system, ${\mathbb{C}}$ is the combination set of \${\mathbb{R}}\backslash\{i\},i\in{\mathbb{R}}$, and $\mathrm{Performance}$ denotes the overall performance of the system on the current task, e.g., classification accuracy or Pass@1. The metric requires ground truth and multi-pass results of the system with different subsets of agents. We use classification accuracy for classification tasks and Pass@1 for code generation tasks. However, its computation cost is still too high when the number of agents grows larger due to its combinatorial complexity.
@@
To examine Shapley Value as an indicator of agent team optimization, we also randomly chose three combinations of three roles out of all 7 roles to assemble DyLAN with three agents and calculated the Shapley Value and the Agent Importance Score in the 3-agent DyLAN on MMLU. In Table 11, we report the correlations between Shapley Values and Agent Importance Scores on the 3-agent experiments. We are curious whether Agent Importance Score is an unsupervised substitution for Shapley Value. So, we calculated the KL divergence and a listwise loss (ListMLE (Xia et al., 2008)) between Agent Importance Scores and Shapley Value. It indeed shows a high correlation between the distributions of the two metrics when at least one agent is in the domain of the question (in column In-Domain).
@@
In summary, while Shapley Value is a self-evident metric for individual contribution, Agent Importance Score emerges as a promising, unsupervised alternative with light computational complexity.
Table 11: Correlation between different metrics for quantifying agents’ contributions in 3-agent DyLAN on MMLU dataset. We compute the KL divergence $D_{\mathrm{KL}}$ and the ListMLE loss $\mathcal{L}_{\mathrm{ListMLE}}$ between Shapley Value and other metrics on each subject and report the average value. The In-Domain column means at least one agent in DyLAN matches the subject according to Appendix C.1, and Off-Domain means none of agents matches the subject.
@@
Importance Score $\mathbf{0.229\times 10^{-3}}$ $\mathbf{0.686}$ $0.347\times 10^{-3}$ $0.693$
@@
Figure 5: A case of DyLAN solving code generation task. Different agents are recruited to write code and give feedback. At the time steps $t=2,5$, we ask judges to provide code reviews. The result grows better layer by layer regarding correctness, efficiency, and readability. Different directions of implementation are delivered forward in implicit multiple paths. We ignore the peer rating scores in responses of agents for computing Agent Importance Scores.
D.6 Case Study
@@
In Figure 5 and Figure 6, we demonstrate the cases of DyLAN on the code generation and the general reasoning tasks, respectively. First, we notice that the interaction architecture is different between figures, exhibiting dynamic architecture of DyLAN on different queries. The former gives answer at $t=4$, the latter at $t=2$. We also notice that the answer is gradually growing better along the temporal axis. In Figure 6, only the agent with “Mathematician” role gives a correct answer and then convince other agents, suggesting that proper agents are critical to task performance and our agent team optimization method is effective. Although there are hallucinations when agents respond to the query or evaluate predecessors’ responses, the overall solution is correct, and the distribution of Agent Importance Score is reasonable. Also, instructing LLM agents to rate scores on predecessors hints them to reflect on predecessors’ responses, which might be helpful to give better answers. Last but not least, agents with different roles lead a diverse conversation and make full use of it, which benefits performance and robustness.
Appendix E Prompting Templates
@@
In DyLAN, agents are assigned roles extracted from an open-source code base [^3] https://github.com/GoGPTAI/ChatGPT-Prompt/blob/main/prompts.csv, relative research projects (Du et al., 2023; Shinn et al., 2023; Zheng et al., 2023), and GPT-4 generation with minor modification. We exhibit the instruction templates of different datasets and the prompts of all roles in Table 12.
@@
Table 12: Instruction and prompting templates used in different datasets and roles.
Role Prompt
MMLU Instruction
@@
MATH Instruction
[…]
HumanEval Instruction
[…]
Mathematician
@@
You are a mathematician. You are good at math games, arithmetic calculation, and long-term planning.
Programmer (MMLU)
@@
You are a programmer. You are good at computer science, engineering, and physics. You have experience in designing and developing computer software and hardware.
Lawyer
You are a lawyer. You are good at law, politics, and history.
Historian
@@
You are a historian. You research and analyze cultural, economic, political, and social events in the past, collect data from primary sources and use it to develop theories about what happened during various periods of history.
Economist
@@
You are an economist. You are good at economics, finance, and business. You have experience on understanding charts while interpreting the macroeconomic environment prevailing across world economies.
Psychologist
@@
You are a psychologist. You are good at psychology, sociology, and philosophy. You give people scientific suggestions that will make them feel better.
Doctor
@@
You are a doctor and come up with creative treatments for illnesses or diseases. You are able to recommend conventional medicines, herbal remedies and other natural alternatives. You also consider the patient’s age, lifestyle and medical history when providing your recommendations.
Python Assistant
@@
You are a Python writing assistant, an AI that only responds with python code, NOT ENGLISH. You will be given a function signature and its docstring by the user. Write your full implementation (restate the function signature).
Algorithm Developer
@@
You are an algorithm developer. You are good at developing and utilizing algorithms to solve problems. You must respond with python code, no free-flowing text (unless in a comment). You will be given a function signature and its docstring by the user. Write your full implementation following the format (restate the function signature).
Computer Scientist
@@
You are a computer scientist. You are good at writing high performance code and recognizing corner cases while solve real problems. You must respond with python code, no free-flowing text (unless in a comment). You will be given a function signature and its docstring by the user. Write your full implementation following the format (restate the function signature).
Programmer (HumanEval)
@@
You are an intelligent programmer. You must complete the python function given to you by the user. And you must follow the format they present when giving your answer! You can only respond with comments and actual code, no free-flowing text (unless in a comment).
Coding Artist
@@
You are a coding artist. You write Python code that is not only functional but also aesthetically pleasing and creative. Your goal is to make the code an art form while maintaining its utility. You will be given a function signature and its docstring by the user. Write your full implementation following the format (restate the function signature).
Software Architect
@@
You are a software architect, skilled in designing and structuring code for scalability, maintainability, and robustness. Your responses should focus on best practices in software design. You will be given a function signature and its docstring by the user. Write your full implementation following the format (restate the function signature).
Unit Tester
@@
You are an AI coding assistant that can write unique, diverse, and intuitive unit tests for functions given the signature and docstring.
Syntax Checker
Null
Code Reflector
@@
You are a Python writing assistant. You will be given a series of function implementations of the same function signature. Write a few sentences to explain whether and why the implementations are wrong. These comments will be used as a hint and your goal is to write your thoughts on the n-th previous implementation after [reflection n].
Debugger
@@
You are a debugger, specialized in finding and fixing bugs in Python code. You will be given a function implementation with a bug in it. Your goal is to identify the bug and provide a corrected implementation. Include comments to explain what was wrong and how it was fixed.
Quality Manager
@@
You are a quality manager, ensuring that the code meets high standards in terms of readability, efficiency, and accuracy. You will be given a function implementation and you need to provide a code review. Comment on its correctness, efficiency, and readability, and suggest improvements if needed.
Ranker
@@
You are a Python writing assistant. You will be given a series of function implementations of the same function signature. You need to choose the best 2 implementations in consideration of correctness, efficiency, and possible corner cases.

### v2

@@ C.3 Human Priors and Agent Importance Scores
Table 11: Subjects on which agents have the top-ranked Agent Importance Score in the experiment with DyLAN of 7 agents on the GR task. Green annotation denotes the fields related to the role from the human perspective, which are annotated manually.
Though most subjects seems to be reasonably aligned with the role of the agent based on human priors (with green annotations), there are some subjects that do not match human priors, e.g., High School Computer Science as the subject that “Doctor” has the highest score. It exhibits the difference between human priors and the evaluation results of agent importance scores on agents with human-made or LLM-generated prompts.
@@
We also compare current agent selection method that is implemented with Agent Importance Score with the implementation with Human Prior Selection on a few subjects in the MMLU and the HumanEval datasets. For Human Prior Selection, we setup GPT-4 mimicking human selecting the agents for collaborations based on the description of the task and role prompts of each agent. We provide prompt templates in Appendix D. As shown in Table 13, the implementation with Agent Importance Score steadily outperforms Human Prior Selection. There are two major reasons: (1) Compared to posterior optimization methods, prior selection may not grasp the actual behaviors of agents, and may not understand which agents are most contributory or helpful to others in the real collaboration process. Thus, in High School Statistics, Clinical Knowledge, and Public Relations subjects in the MMLU dataset, prior selection performs even worse than random selection. (2) Human Prior Selection might struggle to understand tool augmentation without peer ratings from fellow agents. From our observation, “Unit Tester” and “Syntax Checker” were not selected for code generation tasks, which may cause lower performance.
#Agents Optimization College Management High School Clinical Public Overall
Indicator Mathematics Statistics Knowledge Relations
7 (before optimization) 40.0 76.2 65.1 69.8 54.5 63.5 (+0.0)
4 Random Selection 45.0 71.4 67.4 71.7 54.5 64.8 (+1.3)
Human Prior Selection 60.0 80.1 65.1 69.8 54.5 66.7 (+3.2)
Agent Importance Score 65.0 90.5 74.4 75.5 59.1 73.6 (+10.1)
#Agents Optimization Indicator Pass@1 #API Calls
12 (before optimization) 76.2 (+0.0) 23.04
8 Random Selection 75.6 (-0.6) 17.73
Human Prior Selection 78.0 (+1.8) 16.37
Agent Importance Score 82.9 (+6.7) 16.85
Table 13: Detailed performance of different indicators of agent selection on five subjects in GR tasks (top) and the CG task (bottom). The five subjects in GR tasks and other settings are identical to Table 8. The overall accuracy in the top table denotes the accuracy across the five subjects.
C.4 Stability of DyLAN on Temperature
Figure 4: Performance of different methods under low and high temperatures on AR (left) and CG (right) tasks. DyLAN shows better robustness to different temperature and even takes advantage of higher temperature.
We tested a few methods on the AR (with simple CoT prompts) and the CG tasks under both low and high temperatures and repeated each experiment three times when the temperature was not zero. We exhibit the experimental results in Figure 4. From experimental results, we found that DyLAN is more stable on different hyper-parameters.
Experiments show that temperature greatly influences arithmetic reasoning and code generation tasks. In Figure 4, we found that most baseline methods have significant performance drops when temperature increases, but DyLAN shows strong robustness to various temperatures. We surprisingly found that DyLAN gets better results when temperature rises, suggesting it has benefited from diversity instead of being disturbed by low-quality answers of high-temperature agents. The agent team reformation may lead to the higher accuracy by keeping best responses when agents’ replies become more diverse. In conclusion, the collaboration of different roles functions effectively and robustly in the dynamic architecture. Nonetheless, higher temperature requires DyLAN to take more API calls (about +0.98 on average on AR tasks (temperature: $0.2\rightarrow 1.0$)).
C.5 Different Ranking Methods
@@
Table 14: Overall accuracy (%) of DyLAN with different ranking method in the agent team reformation on the GR task. Other settings are identical with Table 5.
@@
We also tested different ranking methods for agent team reformation of DyLAN on the GR task. We tested listwise ranker with our own prompts, pairwise GPT ranker from original LLM-Blender (Jiang et al., 2023), Elo Score from TrueSkill (Herbrich et al., 2006) also implemented with pairwise ranker, and pairwise ranker with Sliding Window algorithm (Qin et al., 2023). In Table 14, we show that different ranking methods have a relatively low impact on performance, probably because of strong discrimination ability of GPT-3.5, but pairwise ranking methods always consume higher computational cost. Thus, we chose a listwise ranker in our implementation of DyLAN.
C.6 Does Agent Importance Score Captures Actual Contributions?
@@
Shapley Value is a widely used supervised metric for evaluating contribution of a single agent in a multi-agent system. Though it is not suitable for unsupervised Team Optimization, by viewing it as a ground-truth metric for measuring individual contributions, we can use it for validating the Agent Importance Score. We implement a simplified algorithm for LLM-powered agent collaboration systems. Given that the collaboration process is symmetric in the formulation of the temporal feed-forward network (Section 3.2), we could reduce the permutation set in the original formula (Lundberg & Lee, 2017) to the combination set:
$S_{i}(\mathcal{R})=\frac{1}{|\mathcal{C}||\mathcal{R}|}\sum_{\mathcal{T}\in\mathcal{C}}\frac{}{}(\mathrm{Performance}(\mathcal{T}\cup\{i\})-\mathrm{Performance}(\mathcal{T})),$ (13)
@@
where $\mathcal{R}$ is the set of agents in the system, $\mathcal{C}$ is the combination set of \$\mathcal{R}\backslash\{i\},i\in\mathcal{R}$, and $\mathrm{Performance}$ denotes the overall performance of the system on the current task, e.g., classification accuracy or Pass@1. The metric requires ground truth and multi-pass results of the system with different subsets of agents. We use classification accuracy for classification tasks and Pass@1 for code generation tasks. However, its computation cost is still too high when the number of agents grows larger due to its combinatorial complexity.
To examine Agent Importance Score as an indicator of agent selection with Shapley Value, we also randomly chose three combinations of three agents out of all 7 candidates to assemble a T-FFN and calculated the Shapley Value and the Agent Importance Score on GR tasks. In the GR task, The roles of candidates in DyLAN match the categories of MMLU in human priors, including “Mathematician” and “Programmer” for STEM, “Lawyer” and “Historian” for Humanities, “Economist” and “Psychologist” in Social Science, and “Doctor” for clinical questions in the “Other” category. During experiment with a T-FFN with at least one agent matches the category of the question, it is called a In-Domain scenario; vice versa.
@@
In Table 15, we report the correlations between Shapley Values and Agent Importance Scores. We are curious whether Agent Importance Score is an unsupervised substitution for Shapley Value. So, we calculated two list-wise metrics for the similarity between distributions: the KL divergence and ListMLE (Xia et al., 2008), between Agent Importance Scores and Shapley Value. It indeed shows a high correlation between the distributions of the two metrics during in-domain scenarios.
@@
In summary, we use Shapley Value as a self-evident metric for measuring individual contribution, showing that Agent Importance Score emerges as a promising, unsupervised alternative with light computational complexity.
@@
Agent Importance Score $\mathbf{0.229\times 10^{-3}}$ $\mathbf{0.686}$ $0.347\times 10^{-3}$ $0.693$
@@
Table 15: Correlation between different metrics for quantifying agents’ contributions in DyLAN on the GR task. We compute the KL divergence ($D_{\mathrm{KL}}$) and the ListMLE loss ($\mathcal{L}_{\mathrm{ListMLE}}$) between Shapley Value and other metrics on each subject and report the average value. The In-Domain column means at least one agent in DyLAN matches the category of the subject according to Section C.6, and Off-Domain means none of agents matches the subject.
Figure 5: A case of DyLAN solving code generation task. Different agents are recruited to write code and give feedback. At the time steps $t=2,5$ code reviewers are asked to provide code reviews. The result grows better layer by layer regarding correctness, efficiency, and readability. Different directions of implementation are delivered forward in implicit multiple paths. We ignore the peer rating scores in multi-round responses for computing Agent Importance Scores due to space limits.
C.7 Case Study
@@
Figure 7: A failed case of DyLAN solving code generation task. Optimization directions provided by Code Reflector are wrong.
@@
In Figure 5 and Figure 6, we demonstrate the cases of DyLAN on the code generation and the general reasoning tasks, respectively. First, we notice that the communication structure is different between figures, exhibiting dynamic architecture of DyLAN on different queries. The former gives answer at $t=4$, the latter at $t=2$. We also notice that the answer is gradually growing better along the temporal axis. In Figure 6, the “Mathematician” agent is selected for the arithmatic query and it gives a correct answer while convincing other agents, agent selection method is effective. We can also observe that the distribution of Agent Importance Score is reasonable. Also, instructing LLM agents to rate scores on predecessors hints them to reflect on predecessors’ responses, which might be helpful to give better answers. Last but not least, agents with different roles lead a diverse conversation and make full use of each one, which benefits performance and robustness.
Last but not least, we provide qualitative analysis on a failure case in Figure 7.
Appendix D Prompt Templates
@@
In DyLAN, agents are assigned roles with prompts extracted from an open-source code base [^3] https://github.com/GoGPTAI/ChatGPT-Prompt/blob/main/prompts.csv, relative research projects (Du et al., 2023; Shinn et al., 2023; Zheng et al., 2023), and generation results of GPT-4-0613, besides manual construction. The prompt of each agent are constructed by concatenation of the role prompt and optional tool descriptions, as the system prompt, and instruction prompts. We exhibit the instruction templates of different datasets, the prompts of all agents, and their sources in LABEL:tab:role-prompts. We annotate the task where each prompt is used in the parenthesis, and the source of each prompt template. We omit the in-context examples of AR tasks from the original dataset of MATH (Hendrycks et al., 2021b) and PHP Zheng et al. (2023), and WebShop from ReAct (Yao et al., 2023).
@@
Table 16: Instruction and prompting templates used in different datasets and agents.
Prompt Content Source
MMLU Instruction (GR) Here is the question: {question}
These are the solutions to the problem from other agents: {responses}
Using the reasoning from other agents as additional advice with critical thinking, can you give an updated answer? Examine your solution and that other agents step by step. Notice that their answers might be all wrong. Put your answer in the form (X) at the end of your response. (X) represents choice (A), (B), (C), or (D).
Manual
MATH Instruction (AR) Follow the given examples and answer the mathematics problem.
{question}
These are the solutions to the problem from other agents: {responses}
Using the reasoning from other agents as additional advice with critical thinking, can you give an updated answer? Examine your solution and that other agents step by step. Notice that their answers might be all wrong.
Manual
HumanEval Instruction (CG) You must complete the python function I give you by rectifying previous implementations. Use the other information as a hint.
Be sure to use the same indentation I specified. Furthermore, you may only write your response in code/comments.
[improved impl]:
```python
{function signature}
Please follow the template by repeating the function signature and complete the new implementation in [improved impl]. If no changes are needed, simply rewrite the implementation in the Python code block.
Retrieved
WebShop Instruction (DM) {history of observatons and actions}
These are the suggested next action from other agents: {actions from other agents}
Using the solutions from other agents as additional advice with critical thinking, can you give an updated action in response to previous observations and actions? Please select the item after searching (click on the product name ‘B0…‘). And when you buy the item (‘click[Buy Now]‘), please make sure you have selected an item and entered it’s description page. Do not search for multiple times. Based on the example and previous actions and observations on the new instruction, give your next action in the format of “Action: search[…]” or “Action: click[…]”.
Retrieved
Human Prior Selection Instruction (-) A few agents will collaborate on the same task query. Please select the optimal composition of the candidate agents based on the description of the task and the agents’ profiles.
- Task: {subject}
- Agents
@@
We want to select k agents/code writers/code reviewers among these candidates. Please write the agent IDs as the following format: [1, 2, 3, 4]. There could be multiple agents with the same ID.
Manual
Ranker Instruction (-) Here is the question: {question}
These are the solutions to the problem from other agents: {responses}
Please choose the best 2 solutions and think step by step. Put your answer in the form like [1,2] or [3,4] at the end of your response.
Manual
@@
Mathematician (GR) You are a mathematician. You are good at math games, arithmetic calculation, and long-term planning. Retrieved
@@
Programmer (GR) You are a programmer. You are good at computer science, engineering, and physics. You have experience in designing and developing computer software and hardware. Retrieved
Lawyer (GR) You are a lawyer. You are good at law, politics, and history. Retrieved
@@
Historian (GR) You are a historian. You research and analyze cultural, economic, political, and social events in the past, collect data from primary sources and use it to develop theories about what happened during various periods of history. Retrieved
@@
Economist (GR) You are an economist. You are good at economics, finance, and business. You have experience on understanding charts while interpreting the macroeconomic environment prevailing across world economies. Retrieved
@@
Psychologist (GR) You are a psychologist. You are good at psychology, sociology, and philosophy. You give people scientific suggestions that will make them feel better. Retrieved
@@
Doctor (GR) You are a doctor and come up with creative treatments for illnesses or diseases. You are able to recommend conventional medicines, herbal remedies and other natural alternatives. You also consider the patient’s age, lifestyle and medical history when providing your recommendations. Retrieved
@@
Python Assistant (CG) You are a Python writing assistant, an AI that only responds with python code, NOT ENGLISH. You will be given a function signature and its docstring by the user. Write your full implementation (restate the function signature). Retrieved
@@
Algorithm Developer (CG) You are an algorithm developer. You are good at developing and utilizing algorithms to solve problems. You must respond with python code, no free-flowing text (unless in a comment). You will be given a function signature and its docstring by the user. Write your full implementation following the format (restate the function signature). Retrieved
@@
Computer Scientist (CG) You are a computer scientist. You are good at writing high performance code and recognizing corner cases while solve real problems. You must respond with python code, no free-flowing text (unless in a comment). You will be given a function signature and its docstring by the user. Write your full implementation following the format (restate the function signature). Retrieved
@@
Programmer (CG) You are an intelligent programmer. You must complete the python function given to you by the user. And you must follow the format they present when giving your answer! You can only respond with comments and actual code, no free-flowing text (unless in a comment). Retrieved
@@
Coding Artist (CG) You are a coding artist. You write Python code that is not only functional but also aesthetically pleasing and creative. Your goal is to make the code an art form while maintaining its utility. You will be given a function signature and its docstring by the user. Write your full implementation following the format (restate the function signature). Generated
@@
Software Architect (CG) You are a software architect, skilled in designing and structuring code for scalability, maintainability, and robustness. Your responses should focus on best practices in software design. You will be given a function signature and its docstring by the user. Write your full implementation following the format (restate the function signature). Generated
@@
Unit Tester (CG) You are an AI coding assistant that can write unique, diverse, and intuitive unit tests for functions given the signature and docstring. Retrieved
Syntax Checker (CG) Null -
@@
Code Reflector (CG) You are a Python writing assistant. You will be given a series of function implementations of the same function signature. Write a few sentences to explain whether and why the implementations are wrong. These comments will be used as a hint and your goal is to write your thoughts on the n-th previous implementation after [reflection n]. Generated
@@
Debugger (CG) You are a debugger, specialized in finding and fixing bugs in Python code. You will be given a function implementation with a bug in it. Your goal is to identify the bug and provide a corrected implementation. Include comments to explain what was wrong and how it was fixed. Generated
@@
Quality Manager (CG) You are a quality manager, ensuring that the code meets high standards in terms of readability, efficiency, and accuracy. You will be given a function implementation and you need to provide a code review. Comment on its correctness, efficiency, and readability, and suggest improvements if needed. Generated
@@
Ranker (CG) You are a Python writing assistant. You will be given a series of function implementations of the same function signature. You need to choose the best 2 implementations in consideration of correctness, efficiency, and possible corner cases. Generated
Search Optimizer (DM) As a Search Optimizer, analyze a user’s vague or broad instruction and suggest a more precise and effective set of keywords or filters to accurately find products that meet their specific requirements. Please focus more on searching actions when giving the next action, especially on providing informative and accurate searching words in “search[…]” Generated
Budget Analyst (DM) As a Budget Analyst, guide a user in setting a realistic budget for their shopping needs, considering product categories, market prices, and personal financial constraints. Analyze whether the product matches the budget one by one. Please avoid searching multiple times. Please focus more on on which product to click in “click[…]” when giving the next action. Generated
Product Explorer (DM) As a Product Explorer, offer guidance on how to effectively compare different products based on their features, prices, and reviews. Advise on key factors to consider for making informed decisions in various product categories. You are preferred to “click[¡ Prev]” or “click[Next ¿]” to go on different pages, and click on the item name to browse its details. Generated
Instruction Analyst (DM) As an Instruction Analyst, evaluate a customer’s stated needs and preferences, and provide a structured approach to ensure the selected product align with these instructions. Do not keep refining searching. You are preferred to check out products that might align with instruction. Otherwise, give the most reasonable action for the next step Generated
Description Reader (DM) As a Description Reader, detail how to read and interpret product descriptions and specifications to match them with a customer’s specific needs, focusing on identifying key features, benefits, and potential drawbacks. You are preferred to click into each product and “click[Description]” to see product details. Generated
Decision Maker (DM) As a Decision Maker, you are more confident to purchase related products without refining searching results. If you believe a product is a suitable choice, proceed to click in the product and persuade other agents to click “click[Buy Now]”. If not, recommend the most logical next step for rapid adaptation for the next product. Generated
Decision Reflector (DM) As a Decision Reflector, provide a framework for customers to critically evaluate their potential purchases, considering their initial requirements, product features, and overall value. Guide them in reflecting on whether a choice truly meets their needs. Please avoid repeated searching. If you think it’s good to buy the product, go “click[Buy Now]”. Otherwise, give the most reasonable action for the next step. Manual
Result Estimater (DM) As a Result Estimator, describe how to predict the potential satisfaction and success of a customer’s purchase decision based on their needs, product choice, and market trends. Offer insights into how these choices may meet their expectations. If you have any suggestions, print them in “think[…]”. Otherwise, give the most reasonable action for the next step. Generated